# Ideal Triangulations and Once-Punctured Surface Bundles

### v1

@@ Abstract.
A well-known result of Walsh states that if $\mathcal{T}^{*}$ is an ideal triangulation of an atoroidal, acylindrical, irreducible, compact 3-manifold with torus boundary components, then every properly embedded, two-sided, incompressible surface $S$ is isotopic to a spun-normal surface unless $S$ is isotopic to a fiber or virtual fiber. Previously it was unknown if for such a 3-manifold an ideal triangulation in which a fiber spun-normalizes exists. We give a proof of existence and give an algorithm to construct the ideal triangulation provided the 3-manifold has a single boundary component.
@@ 1. Introduction
Normal and spun-normal surface theory lies at the heart of algorithmic 3-manifold topology. Since Haken’s foundational work on unknot recognition [18], normal surfaces have played a central role in key algorithmic advances, including Thompson’s [41] and Rubinstein’s [38] solutions to the 3-sphere recognition problem, and the detection of reducible surgeries on knots by Jaco, Rubinstein, and Sedgwick [24]. Spun-normal surfaces have proven highly effective in computational topology. They connect deeply with topological quantum invariants, such as the 3D Index, as computed by Garoufalidis, Hodgson, Hoffman, and Rubinstein [17], and the colored Jones polynomial, studied by Futer, Kalfagianni, and Purcell [16]. Especially relevant to this work is the algorithm of Cooper, Tillmann, and Worden [11], which computes the unit Thurston norm ball for cusped hyperbolic 3-manifolds with torus boundary.
Following the groundbreaking results of Bergeron and Wise [4] in 2012 and Agol [2] in 2013, it is known that every hyperbolic 3-manifold (in our case with torus cusps) is virtually fibered—that is, it admits a finite cover which fibers over the circle (we include fibers in this notion of virtual fibers). The virtual fibers are carried by faces of the Thurston norm ball [42]. Yet, virtual fibers are at issue with spun-normal surface theory. A well known theorem by Walsh [45] illustrates this obstacle: in a triangulation with essential edges, every properly embedded, two-sided, incompressible, $\partial$-incompressible surface that is not a virtual fiber is isotopic to a spun-normal surface. This highlights a gap: virtual fibers typically fall outside the reach of spun-normal surface techniques. Cooper, Tillmann, and Worden address this issue directly in their work (Question 5.1 of [11]), asking whether every once-cusped, fibered hyperbolic 3-manifold admits an ideal triangulation in which the fiber is isotopic to an embedded spun-normal surface. The aim of this paper is to give an affirmative answer to that question and to present an algorithm for constructing such a triangulation.
@@
Section 3 discusses two complementary operations on triangulations: crushing normal surfaces and inflating ideal triangulations. Crushing triangulations along normal surfaces is a technique developed by Jaco and Rubinstein in [21] to reduce the complexity of 3-manifolds. A review of this process is given in Subsection 3.1. Subsection 3.2 covers the inverse process– inflating ideal triangulations–developed by Jaco and Rubinstein in [23]. Particular attention is paid here to the boundary of the inflation triangulations and how the push-off of the boundary intersects individual tetrahedra. In Subsection 3.3, we introduce the notion of short inflations, those having a minimal number of inflated tetrahedra. This minimality plays a central role in our approach and, when combined with the concept of compatible normal surfaces, forms the foundation for the two main technical results of this work:
@@
[3.13] Let $M$ be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty boundary homeomorphic to a torus. Further, let $\mathcal{T}$ be a short inflation with $S$ a compatible normal surface. The ideal triangulation $\mathcal{T}^{*}$ obtained by crushing $\mathcal{T}$ along $\partial M$ has a spun-normal surface $S^{*}$ with normal quadrilaterals completely determined by the normal quadrilaterals of $S$ in $\mathcal{T}$, the lifts of which (relative to the crushing map) are contained in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$. Further $S^{*}$ represents the interior of a properly embedded surface in $\mathring{M}$ isotopic to $\mathring{S}$.
@@
Section 5 gives two examples of the implementation of Theorem 4.1: first for the complement of the trefoil knot, and second, for the complement of the (-2,3,7)-pretzel knot. In both cases, the ideal triangulations constructed are, to the best of our knowledge, the first in the literature with the property that the fiber ($S_{1,1}$ and $S_{5,1}$ resp.) is isotopic to an embedded spun-normal surface. The section concludes with a set of open questions aimed at extending the study of such triangulations.
@@ Acknowledgments
The author would like to thank William Jaco for initiating this investigation. The author would like to express their gratitude to Saul Schleimer for key references and encouraging the focus on vertex solutions over fundamental solutions. Finally, the author would like to thank Ian Agol and Joel Hass for stimulating conversations leading to some of the specific examples in Section 5.
@@ 2. Triangulations and Normal Surface Theory
We adopt the notation used in [21] for triangulations and ideal triangulations. Let $\widetilde{\mathbf{\Delta}}=\sqcup_{i}\Delta_{i}$ be a collection of pairwise disjoint tetrahedra, $\mathbf{\Phi}$ a collection of orientation-reversing face identifications, or glueings. The identification space $X=\widetilde{\mathbf{\Delta}}/\mathbf{\Phi}$ is a 3-manifold at each point except possibly at a collection of vertices, $\{v_{i}\}$. If $X=M$ is a manifold, we say $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ is a triangulation of $M$. If $X$ is not a manifold, $X\backslash\{v_{i}\}$ is homeomorphic to the interior of a compact 3-manifold with boundary, $M$, and we say $\mathcal{T}^{*}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ is an ideal triangulation of $M$ and we refer to the tetrahedra of $\mathcal{T}^{*}$ as ideal tetrahedra. We denote the image of the $k$-subsimplices of $\widetilde{\mathbf{\Delta}}$ under the quotient map by $\mathcal{T}^{(k)}$ (${\mathcal{T}^{*}}^{(k)}$).
Throughout this paper, $M$, will be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty connected boundary homeomorphic to a torus. By work of Jaco and Rubinstein [21], we can assume that an ideal triangulation, $\mathcal{T}^{*}$ of $M$, has a single vertex, $v^{*}$, referred to as the ideal vertex. The frontier of a regular neighborhood $N(v^{*})$ we call the ideal boundary of $\mathcal{T}^{*}$. For a triangulation $\mathcal{T}$ of $M$ it is necessary for some tetrahedra to have unglued faces. These unglued faces triangulate $\partial M$, we refer to this decomposition of $\partial M$ as the material boundary of $\mathcal{T}$.
In our construction, the tetrahedra of $\widetilde{\mathbf{\Delta}}$ are not necessarily embedded in $X$, but their interiors are always embedded. For each tetrahedron, $\Delta$ in $X$, there is exactly one tetrahedron, $\widetilde{\Delta}$ in $\widetilde{\mathbf{\Delta}}$, that projects to $\Delta$, we refer to $\widetilde{\Delta}$ as the lift of $\Delta$. A face, $\sigma$ in $X$ lifts to two faces in $\widetilde{\mathbf{\Delta}}$ (not necessarily on distinct tetrahedra) if and only if the interior of $\sigma$ is in the interior of $X$. Otherwise $\sigma$ lifts to exactly one face in $\widetilde{\mathbf{\Delta}}$ and $\sigma$ is in the material boundary of $\mathcal{T}$.
For the 3-manifolds of interest in this work, ideal-triangulations can be constructed using numerous methods. Bing [5] gives a method to construct a triangulation of $M$ having material boundary. This material boundary can be collapsed to an ideal vertex using the methods of Jaco and Rubinstein [21]. In [33], Lackenby gives an algorithm to construct a “taut” ideal triangulation, the faces of which carry all properly embedded, $\chi_{-}$-minimizing incompressible surfaces.
@@ 2.1. Normal and Spun-Normal Surface Theory
Let $\mathcal{T}$ be a triangulation of $M$. Let $S$ be a properly embedded surface in $M$ which is transverse to $\mathcal{T}^{(2)}$, we say that $S$ is normal (normal in $\mathcal{T}$) if for each tetrahedron $\Delta$ of $\mathcal{T}$ the intersection of $S$ with $\Delta$ lifts to a finite collection of normal triangles and normal quadrilaterals in $\widetilde{\Delta}$. An isotopy of $M$ is called a normal isotopy (with respect to $\mathcal{T}$) if it is invariant on each simplex of $\mathcal{T}$. A triangulation $\mathcal{T}$ is said to have normal boundary if the frontier of a regular neighborhood of $\partial M$ is normally isotopic to a normal surface in $\mathcal{T}$. We call said normal surface a boundary-linking surface. It is not necessary in general for a triangulation with material boundary to have normal boundary, but it will be so for the triangulations constructed later in this work.
@@
For a given triangulation, $\mathcal{T}$, with $t=|\mathcal{T}^{(3)}|$ tetrahedra there are $7t$ normal isotopy classes of normal discs. Given an ordering of these normal disc types $d_{1},d_{2},\dots d_{7t}$ we can represent a given normal surface $S$ by the vector $(x_{1},x_{2},\dots,x_{7t})\in\mathbb{R}^{7t}_{\geq 0}$ where $x_{i}$ is the number of normal discs in the induced cell-decomposition of $S$ of type $d_{i}$. Suppose $S$ meets a face $\sigma$ of $\mathcal{T}$ in a normal arc $\gamma$. For each of $\widetilde{\Delta}_{i},\widetilde{\Delta}_{j}$ (possible $i=j$) identified at $\sigma$, $\gamma$ determines one normal triangle type and one normal quadrilateral type. Let $d_{i},d_{j},d_{k},d_{l}$ be those types, then the face identification of $\mathcal{T}$ determines a matching equations: $x_{i}+x_{j}=x_{k}+x_{l}$. For each face, $\sigma$ in $\mathcal{T}$ that lifts to two faces of some $\widetilde{\Delta}_{i},\widetilde{\Delta}_{j}\in\widetilde{\mathbf{\Delta}}$, there are three such equations determined by the three normal arc classes in $\sigma$. Non-negative integer solutions to this system give a parametrization of the normal isotopy classes of normal surfaces in $\mathcal{T}$. We add the restriction that the solutions to the matching equations have non-negative coordinates and we get a cone in the positive-orthant of $\mathbb{R}^{7t}_{\geq 0}$. We refer to this cone as the solution space, $\mathcal{S}_{\mathcal{T}}$. We refer to the integer lattice points of $\mathcal{S}_{\mathcal{T}}$ as solutions. For a normal surface, $S$, we refer to both the embedding and the normal isotopy class as simply $S$. If in addition to the gluing equations we require $\sum x_{i}=1$, we determine a compact, convex linear cell. The rational points in this cell correspond to the projective classes of points in $\mathcal{S}_{\mathcal{T}}$. We refer to this compact, convex linear cell as the projective solution space, $\mathcal{P}_{\mathcal{T}}$. A rational point in $\mathcal{P}_{\mathcal{T}}$ is an admissible solution if corresponding to each tetrahedra there is at most one of the normal quadrilateral types whose corresponding coordinate is non-zero. We denote the projective class of a normal surface $S$ by $\vec{S}\in\mathcal{P}_{\mathcal{T}}$. Every admissible solution in $\mathcal{P}_{\mathcal{T}}$ is the projective class of some normal surface in $\mathcal{T}$.
@@
The carrier, $C_{S}$, of a normal surface $S$ is the unique minimal face of $\mathcal{P}_{\mathcal{T}}$ that contains $\vec{S}$. A normal surface $F$ is said to be carried by $C_{S}$ if $\vec{F}\in C_{S}$. We have that $F$ is carried by $C_{S}$ if and only if there exists a normal surface $G$ such that $nS=F+G$ for some positive integer $n$. Every rational point of $C_{S}$ is an admissible solution. In particular, a vertex $\vec{v}$ of $C_{S}$ (more generally a vertex of $\mathcal{P}_{\mathcal{T}}$) is a rational point and must be admissible as each coordinate corresponding to a normal disc type not in $S$ is zero. We say a normal surface, $S$, is a vertex solution if its parametrization is $k\vec{v}$ where $k$ is the smallest integer for which $k\vec{v}$ has non-negative integer coordinates.
@@
In [43], Tollefson showed that for a triangulation $\mathcal{T}$ the admissible solutions are completely determined by their normal quadrilaterals. Kang [28] extended this result to show for ideal triangulations, both normal and spun-normal surfaces are completely determined by their normal quadrilaterals. In both instances, the notation will be as follows: given a (ideal) triangulation with $t$ tetrahedra, there are $3t$ normal isotopy classes of normal quadrilaterals. Fix an ordering of these normal quadrilateral types $d_{1}^{Q},d_{2}^{Q},\dots,d_{3t}^{Q}$. There is a unique representation of a (spun-)normal surface by the vector $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$. Consider an edge $e_{k}=\langle ab\rangle$ with positive orientation from $a$ to $b$ in ($\mathcal{T}^{*}$) $\mathcal{T}$. Let $d^{Q}$ be a normal quadrilateral type in $\widetilde{\Delta}$ and let $v_{j}$ be a corner of a normal quadrilateral of type $d^{Q}$ incident to $\hat{e}_{k}=\langle\hat{a}\hat{b}\rangle$ an edge of $\widetilde{\Delta}$ in the lift of $e_{k}$. Each disc of type $d^{Q}$ has within its boundary, a normal arc separating vertex $\hat{a}$ in a face of $\widetilde{\Delta}$ containing $\hat{e}_{k}$ and a normal arc separating vertex $\hat{b}$ in the other face containing $\hat{e}_{k}$. If for the disc type $d^{Q}$ a right handed twist around $\hat{e}_{k}$ moves from the face of $\widetilde{\Delta}$ containing the normal arc of $d^{Q}$ separating $\hat{b}$ to the face containing the normal arc separating $\hat{a}$, we define the sign of the corner $v_{j}$ with respect to the edge $e_{k}$ to be $\delta_{k,j}=+1$. If a right hand twist around $\hat{e}_{k}$ moves from the arc $\hat{a}$ to $\hat{b}$, then we set $\delta_{k,j}=-1$. For every corner $v_{j}$ that is incident to no lift $\hat{e}_{k}$ the sign is defined to be $\delta_{k,j}=0$. For a normal quadrilateral of type $d^{Q}$ we define the sense $\epsilon_{k}$ relative to the edge $e_{k}$ to be the total of the signs of each of the four corners of a quadrilateral of the given type. Consider for a (spun-)normal surface $S$, all normal quadrilateral types incident to some edge $e_{k}$ in (ideal) triangulation $\mathcal{T}$. The number of normal quadrilaterals in the cell decomposition of $S$ of positive sense must be equal to the number of those of negative sense. The parametrization $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$ is a solution to the linear system of equations, the Q-matching equations with one equation for each edge $e_{k}$ in $\mathcal{T}$:

### v2

@@ Abstract.
A well-known result of Walsh states that if $\mathcal{T}^{*}$ is an ideal triangulation of an atoroidal, acylindrical, irreducible, compact 3-manifold with torus boundary components and $\mathcal{T}^{*}$ has essential edges, then every properly embedded, two-sided, incompressible surface $S$ is isotopic to a spun-normal surface in $\mathcal{T}^{*}$ unless $S$ is isotopic to a fiber or virtual fiber. For a given manifold $M$ that fibers over $S^{1}$, it was previously unknown whether there exists an ideal triangulation in which the fiber appears as a spun-normal surface. We prove that such a triangulation exists and give an algorithm to construct the ideal triangulation provided $M$ has a single boundary component.
@@ 1. Introduction
Normal and spun-normal surface theory lies at the heart of algorithmic 3-manifold topology. Since Haken’s foundational work on unknot recognition [18], normal surfaces have played a central role in key algorithmic advances, including Thompson’s [40] and Rubinstein’s [37] solutions to the 3-sphere recognition problem, and the detection of reducible surgeries on knots by Jaco, Rubinstein, and Sedgwick [24]. Spun-normal surfaces have proven highly effective in computational topology. They connect deeply with topological quantum invariants, such as the 3D Index, as computed by Garoufalidis, Hodgson, Hoffman, and Rubinstein [17], and the colored Jones polynomial, studied by Futer, Kalfagianni, and Purcell [16]. Especially relevant to this work is the algorithm of Cooper, Tillmann, and Worden [11], which utilizes embedded and immersed spun-normal surfaces to compute the unit Thurston norm ball for cusped hyperbolic 3-manifolds with torus boundary.
Following the groundbreaking results of Bergeron and Wise [4] in 2012 and Agol [2] in 2013, it is known that every hyperbolic 3-manifold (in our case with torus cusps) is virtually fibered—that is, it admits a finite cover which fibers over the circle (we include fibers in this notion of virtual fibers). The virtual fibers are carried by faces of the Thurston norm ball [41]. Yet virtual fibers are at issue with spun-normal surface theory. A well-known theorem by Walsh [44] illustrates this obstacle: in a triangulation with essential edges, every properly embedded, two-sided, incompressible, $\partial$-incompressible surface that is not a virtual fiber is isotopic to a spun-normal surface. This highlights a gap: virtual fibers typically fall outside the reach of spun-normal surface techniques. Cooper, Tillmann, and Worden address this issue directly in their work (Question 5.1 of [11]), asking whether every fibered, once-cusped, hyperbolic 3-manifold admits an ideal triangulation in which the fiber is isotopic to an embedded spun-normal surface. The aim of this paper is to give an affirmative answer to that question and to present an algorithm for constructing such a triangulation.
@@
Section 3 discusses two complementary operations on triangulations: crushing along normal surfaces and inflating ideal triangulations. Crushing triangulations along normal surfaces is a technique developed by Jaco and Rubinstein in [21] to reduce the complexity of 3-manifolds. A review of this process is given in Subsection 3.1. Subsection 3.2 covers the inverse process–inflating ideal triangulations–developed by Jaco and Rubinstein in [23]. Particular attention is paid here to the boundary of the inflation triangulations and how the push-off of the boundary intersects individual tetrahedra. In Subsection 3.3, we introduce the notion of short inflations, those having a minimal number of inflated tetrahedra. This minimality plays a central role in our approach and, when combined with the concept of compatible normal surfaces, forms the foundation for the two main technical results of this work:
@@
[3.13] Let $M$ be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty connected boundary homeomorphic to a torus. Let $\mathcal{T}$ be a short inflation with $S$ a compatible normal surface. The ideal triangulation $\mathcal{T}^{*}$ obtained by crushing $\mathcal{T}$ along $\partial M$ has a spun-normal surface $S^{*}$ with normal quadrilaterals completely determined by the normal quadrilaterals of $S$ in $\mathcal{T}$, the lifts of which (relative to the crushing map) are contained in $\widetilde{\mathbf{\Delta}}^{*}\subset\widetilde{\mathbf{\Delta}}$. Further, $S^{*}$ represents the interior of a properly embedded surface in $\mathring{M}$ isotopic to $\mathring{S}$.
@@
Section 5 gives two examples of the implementation of Theorem 4.1: first for the complement of the trefoil knot, and second, for the complement of the (-2,3,7)-pretzel knot. In both cases, the ideal triangulations constructed are, to the best of the author’s knowledge, the first in the literature with the property that the fiber ($S_{1,1}$ and $S_{5,1}$ resp.) is isotopic to an embedded spun-normal surface. The section concludes with a set of open questions aimed at extending the study of such triangulations.
@@ Acknowledgments
This work began as the author’s PhD dissertation, supervised by Neil Hoffmann. The author would like to thank him for his patience and continued support. The author would like to thank William Jaco for initiating this investigation. The author would like to express their gratitude to Saul Schleimer for key references and encouraging the focus on vertex solutions over fundamental solutions. The author gratefully acknowledges that part of the work and code used to generate examples in this paper were developed with support from Henry Segerman’s NSF grant (DMS-1708239, Three- and Four-Dimensional Triangulations and Mathematical Visualization, 2017–2021) during a summer research assistantship at Oklahoma State University. Finally, the author would like to thank Ian Agol and Joel Hass for stimulating conversations leading to some of the specific examples in Section 5.
@@ 2. Triangulations and Normal Surface Theory
We adopt the notation used in [21] for triangulations and ideal triangulations. Let $\widetilde{\mathbf{\Delta}}=\{\widetilde{\Delta}_{i}\}_{i}$ be a collection of pairwise disjoint tetrahedra, $\mathbf{\Phi}$ a collection of orientation-reversing face identifications, or glueings. The identification space $X=\widetilde{\mathbf{\Delta}}/\mathbf{\Phi}$ is a 3-manifold at each point except possibly at a collection of vertices, $\{v_{i}\}$. If $X=M$ is a manifold, we say $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ is a triangulation of $M$. If $X$ is not a manifold, $X\backslash\{v_{i}\}$ is homeomorphic to the interior of a compact 3-manifold with boundary, $M$, and we say $\mathcal{T}^{*}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ is an ideal triangulation of $M$ and we refer to the tetrahedra of $\mathcal{T}^{*}$ as ideal tetrahedra. We denote the image of the $k$-subsimplices of $\widetilde{\mathbf{\Delta}}$ under the quotient map by $\mathcal{T}^{(k)}$ (or ${\mathcal{T}^{*}}^{(k)}$).
Throughout this paper, $M$ will be an orientable, compact, irreducible, $\partial$-irreducible 3-manifold with non-empty connected boundary homeomorphic to a torus. By work of Jaco and Rubinstein [21], we can assume that an ideal triangulation, $\mathcal{T}^{*}$ of $M$, has a single vertex, $v^{*}$, referred to as the ideal vertex. The frontier of a regular neighborhood $N(v^{*})$ we call the ideal boundary of $\mathcal{T}^{*}$. For a triangulation $\mathcal{T}$ of $M$ it is necessary for some tetrahedra to have unglued faces. These unglued faces triangulate $\partial M$, we refer to this decomposition of $\partial M$ as the material boundary of $\mathcal{T}$.
In our construction, the tetrahedra of $\widetilde{\mathbf{\Delta}}$ are not necessarily embedded in $X$, but their interiors are always embedded. For each tetrahedron $\Delta$ in $X$, there is exactly one tetrahedron $\widetilde{\Delta}\in\widetilde{\mathbf{\Delta}}$, that projects to $\Delta$, we refer to $\widetilde{\Delta}$ as the lift of $\Delta$. A face, $\sigma$ in $X$ lifts to two faces in $\widetilde{\mathbf{\Delta}}$ (not necessarily on distinct tetrahedra) if and only if the interior of $\sigma$ is in the interior of $X$. Otherwise $\sigma$ lifts to exactly one face in $\widetilde{\mathbf{\Delta}}$ and $\sigma$ is in the material boundary of $\mathcal{T}$.
For the 3-manifolds of interest in this work, ideal-triangulations can be constructed using numerous methods. Bing [5] gives a method to construct a triangulation of $M$ having material boundary. This material boundary can be collapsed to an ideal vertex using the methods of Jaco and Rubinstein [21]. In [33], Lackenby gives an algorithm to construct a taut ideal triangulation, the faces of which carry all properly embedded, $\chi_{-}$-minimizing incompressible surfaces.
@@ 2.1. Normal and Spun-Normal Surface Theory
Let $\mathcal{T}$ be a triangulation of $M$. Let $S$ be a properly embedded surface in $M$ which is transverse to $\mathcal{T}^{(2)}$, we say that $S$ is normal (normal in $\mathcal{T}$) if for each tetrahedron $\Delta$ of $\mathcal{T}$, the intersection of $S$ with $\Delta$ lifts to a finite collection of normal triangles and normal quadrilaterals in $\widetilde{\Delta}$. An isotopy of $M$ is called a normal isotopy (with respect to $\mathcal{T}$) if it is invariant on each simplex of $\mathcal{T}$. A triangulation $\mathcal{T}$ is said to have normal boundary if the frontier of a regular neighborhood of $\partial M$ is normally isotopic to a normal surface in $\mathcal{T}$. We call said normal surface a boundary-linking surface. It is not necessary in general for a triangulation with material boundary to have normal boundary, but it will be so for the triangulations constructed later in this work.
@@
For a given triangulation, $\mathcal{T}$, with $t=|\mathcal{T}^{(3)}|$ tetrahedra there are $7t$ normal isotopy classes of normal discs. Given an ordering of these normal disc types $d_{1},d_{2},\dots d_{7t}$ we can represent a given normal surface $S$ by the vector $(x_{1},x_{2},\dots,x_{7t})\in\mathbb{R}^{7t}_{\geq 0}$ where $x_{i}$ is the number of normal discs in the induced cell-decomposition of $S$ of type $d_{i}$. Suppose $S$ meets a face $\sigma$ of $\mathcal{T}$ in a normal arc $\gamma$. For each of $\widetilde{\Delta}_{i},\widetilde{\Delta}_{j}$ (possible $i=j$) identified at $\sigma$, $\gamma$ determines one normal triangle type and one normal quadrilateral type. Let $d_{i},d_{j},d_{k},d_{l}$ be those types, then the face identification of $\mathcal{T}$ determines a matching equations: $x_{i}+x_{j}=x_{k}+x_{l}$. For each face $\sigma$ in $\mathcal{T}$ that lifts to two faces of some $\widetilde{\Delta}_{i},\widetilde{\Delta}_{j}\in\widetilde{\mathbf{\Delta}}$, there are three such equations determined by the three normal arc classes in $\sigma$. Non-negative integer solutions to this system give a parametrization of the normal isotopy classes of normal surfaces in $\mathcal{T}$. We add the restriction that the solutions to the matching equations have non-negative coordinates and we get a cone in the positive-orthant of $\mathbb{R}^{7t}_{\geq 0}$. We refer to this cone as the solution space, $\mathcal{S}_{\mathcal{T}}$. We refer to the integer lattice points of $\mathcal{S}_{\mathcal{T}}$ as solutions. For a normal surface $S$, we refer to both the embedding and the normal isotopy class as simply $S$. If in addition to the gluing equations we require $\sum x_{i}=1$, we determine a compact, convex linear cell. The rational points in this cell correspond to the projective classes of points in $\mathcal{S}_{\mathcal{T}}$. We refer to this compact, convex linear cell as the projective solution space denoted $\mathcal{P}_{\mathcal{T}}$. A rational point in $\mathcal{P}_{\mathcal{T}}$ is an admissible solution if corresponding to each tetrahedra there is at most one of the normal quadrilateral types whose corresponding coordinate is non-zero. We denote the projective class of a normal surface $S$ by $\vec{S}\in\mathcal{P}_{\mathcal{T}}$. Every admissible solution in $\mathcal{P}_{\mathcal{T}}$ is the projective class of some normal surface in $\mathcal{T}$.
@@
The carrier $C_{S}$ of a normal surface $S$ is the unique minimal face of $\mathcal{P}_{\mathcal{T}}$ that contains $\vec{S}$. A normal surface $F$ is said to be carried by $C_{S}$ if $\vec{F}\in C_{S}$. We have that $F$ is carried by $C_{S}$ if and only if there exists a normal surface $G$ such that $nS=F+G$ for some positive integer $n$. Every rational point of $C_{S}$ is an admissible solution. In particular, a vertex $\vec{v}$ of $C_{S}$ (more generally a vertex of $\mathcal{P}_{\mathcal{T}}$) is a rational point and must be admissible as each coordinate corresponding to a normal disc type not in $S$ is zero. We say a normal surface is a vertex solution if its parametrization is $k\vec{v}$ where $k$ is the smallest integer for which $k\vec{v}$ has non-negative integer coordinates.
@@
In [42], Tollefson showed that for a triangulation $\mathcal{T}$ the admissible solutions are completely determined by their normal quadrilaterals. Kang [28] extended this result to show for ideal triangulations, both normal and spun-normal surfaces are completely determined by their normal quadrilaterals. In both instances, the notation will be as follows: given a (ideal) triangulation with $t$ tetrahedra, there are $3t$ normal isotopy classes of normal quadrilaterals. Fix an ordering of these normal quadrilateral types $d_{1}^{Q},d_{2}^{Q},\dots,d_{3t}^{Q}$. There is a unique representation of a (spun-)normal surface by the vector $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$. Consider an edge $e_{k}=\langle ab\rangle$ with positive orientation from $a$ to $b$ in ($\mathcal{T}^{*}$) $\mathcal{T}$. Let $d^{Q}$ be a normal quadrilateral type in $\widetilde{\Delta}$ and let $v_{j}$ be a corner of a normal quadrilateral of type $d^{Q}$ incident to $\hat{e}_{k}=\langle\hat{a}\hat{b}\rangle$ an edge of $\widetilde{\Delta}$ in the lift of $e_{k}$. Each disc of type $d^{Q}$ has within its boundary, a normal arc separating vertex $\hat{a}$ in a face of $\widetilde{\Delta}$ containing $\hat{e}_{k}$ and a normal arc separating vertex $\hat{b}$ in the other face containing $\hat{e}_{k}$. If for the disc type $d^{Q}$ a right handed twist around $\hat{e}_{k}$ moves from the face of $\widetilde{\Delta}$ containing the normal arc of $d^{Q}$ separating $\hat{b}$ to the face containing the normal arc separating $\hat{a}$, we define the sign of the corner $v_{j}$ with respect to the edge $e_{k}$ to be $\delta_{k,j}=+1$. If a right hand twist around $\hat{e}_{k}$ moves from the arc $\hat{a}$ to $\hat{b}$, then we set $\delta_{k,j}=-1$. For every corner $v_{j}$ that is incident to no lift $\hat{e}_{k}$ the sign is defined to be $\delta_{k,j}=0$. For a normal quadrilateral of type $d^{Q}$ we define the sense $\epsilon_{k}$ relative to the edge $e_{k}$ to be the total of the signs of each of the four corners of a quadrilateral of the given type. Consider for a (spun-)normal surface $S$ all normal quadrilateral types incident to some edge $e_{k}$ in (ideal) triangulation $\mathcal{T}$. The number of normal quadrilaterals in the cell decomposition of $S$ of positive sense must be equal to the number of those of negative sense. The parametrization $S_{Q}=(q_{1},q_{2},\dots,q_{3t})\in\mathbb{R}^{3t}_{\geq 0}$ is a solution to the linear system of equations, the Q-matching equations with one equation for each edge $e_{k}$ in $\mathcal{T}$: